OpenStudy (anonymous):

find the inverse alegrbraically: f(x)=4(^x)+8

OpenStudy (johnweldon1993):

well how do you find the inverse? switch the x's and y's so y = 4^x + 8 becomes x = 4^y + 8 solve for y again

OpenStudy (anonymous):

How do I solve for y?

OpenStudy (anonymous):

subtract 8, then take the log base 4

OpenStudy (anonymous):

x-8=log4?

OpenStudy (anonymous):

you are solving for \(y\) right?

OpenStudy (anonymous):

im trying to find the inverse so i guess so

OpenStudy (anonymous):

you have \[x = 4^y + 8\] subtract 8 get \[x-8=4^y\]

OpenStudy (anonymous):

then take the log to solve for \(y\) and get \[\log_4(x-8)=y\]

OpenStudy (anonymous):

So that is the complete answer?

OpenStudy (anonymous):

as they say on tv, the "final answer"

OpenStudy (anonymous):

thank you

OpenStudy (johnweldon1993):

thanks for taking that one over @satellite I was pre-occupied :)